# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

### v1

@@ Abstract
This article is concerned with networked estimation in a system composed of a source that is observed by a sensor, a remote monitor that needs to estimate the state of the source in real time, and a communication channel that connects the source to the monitor. The source is a partially observable dynamical process, and the communication channel is a packet-erasure channel with feedback. Our main objective is to obtain the fundamental performance limits of the underlying networked system in the sense of a causal tradeoff between the packet rate and the mean square error when both forward and backward channels are unreliable. We characterize an optimal coding policy profile consisting of a scheduling policy for the encoder and an estimation policy for the decoder. We complement our theoretical results with a numerical analysis, and compare the performance limits of the networked system in different communication regimes.
Index Terms:
communication channels, feedback, optimal policies, packet loss, state estimators, unreliability.
@@ I Introduction
Cyber-physical systems represent complex real-time systems that intricately integrate computational algorithms with dynamical processes, thereby facilitating seamless interactions between the digital and physical realms[1]. These systems harness computation, communication, and control to optimize efficiency, adaptability, and automation across diverse domains, including smart factories, smart healthcare, and smart transportation. A significant challenge within cyber-physical systems arises from their inherent dynamism, which requires continuous status updating to capture the latest information about the physical environment. As a step towards addressing this challenge, in this article, we develop a framework for networked estimation in a system composed of a source that is observed by a sensor, a remote monitor that needs to estimate the state of the source in real time, and a communication channel that connects the source to the monitor. The source is a partially observable dynamical process, and the communication channel is a packet-erasure channel with feedback. This framework is motivated by cyber-physical systems operating over digital networks, where the quality of decision making tightly depends on the availability of sensory observations transmitted as data packets, while packet loss appears as an inevitable phenomenon due to the real-time nature of the operation. Our main objective in this article is to obtain the fundamental performance limits of the underlying networked system in the sense of a causal tradeoff between the packet rate and the mean square error when both forward and backward channels are unreliable. To that end, we seek for an optimal coding policy profile that determines the policies of an encoder and a decoder, which are collocated with the source and the monitor, respectively.
@@ I-A Related Work
It is known that reliable communication close to the capacity limit can be attained with error correction subject to infinite delay or with persistent retransmission based on feedback. In networked estimation, where data is real-time, any delay more than a certain threshold is typically intolerable. Moreover, retransmission of stale information is unfavourable when fresh information can be transmitted with the same success rate instead. For these reasons, networked estimation is often conducted in an unreliable communication regime. Previous research has already recognized the severe effects of packet loss, as a consequence of communication unreliability, on stability of Kalman filtering [2, 3, 4, 5, 6, 7, 8, 9, 10]. In a seminal work, Sinopoli et al. [2] studied mean-square stability of Kalman filtering over an independent and identically distributed (i.i.d.) channel, and proved that there exists a critical point for the packet error rate above which the expected estimation error covariance is unbounded. To capture the temporal correlation of wireless channels, Wu et al. [6] addressed peak-covariance stability of Kalman filtering over a Gilbert-Elliott channel, and proved that there exists a critical region defined by the recovery and failure rates outside which the expected prediction error covariance is unbounded. In addition, to take into account the time variation of the strengths of wireless channels, Quevedo et al. [9] investigated mean-square stability of Kalman filtering over a fading channel with correlated gains, and established a sufficient condition that ensures the exponential boundedness of the expected estimation error covariance. Note that, the above works assume that the outputs of the source are transmitted in a periodic manner, and seek to derive fundamental stability conditions regardless of the existence of feedback.
@@
To determine the fundamental performance limits of networked estimation, however, one requires to make a tradeoff between communication and estimation costs. Various classes of such a tradeoff have been explored in the literature [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In particular, the tradeoff between the packet rate and the mean square error, in which a signal-dependent scheduling policy and an estimation policy need to be found, was considered in [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 30, 31]. In this class, which is of main interest to our study, Lipsa and Martins [12, 13] analysed the estimation of a scalar Gauss–Markov process over a lossless channel and an i.i.d. channel with reliable feedback, and proved based on majorisation theory that the optimal scheduling policy is symmetric threshold and the optimal estimation policy is linear. Molin and Hirche [14] investigated the estimation of a scalar Markov process with symmetric noise distribution over a lossless channel, and devised an iterative algorithm that produces policies similar to those found in [12]. Chakravorty and Mahajan [15, 16, 17] studied the estimation of a scalar autoregressive Markov process with symmetric noise distribution over a lossless channel, an i.i.d. channel with reliable feedback, as well as a Gilbert–Elliott channel with reliable feedback, and proved based on renewal theory that the optimal scheduling policy is symmetric threshold and the optimal estimation policy is linear. Moreover, Guo and Kostina [19] addressed the estimation of the scalar Ornstein–Uhlenbeck process over a lossless channel, and based on majorisation theory and real induction proved that the optimal scheduling policy remains symmetric threshold and the optimal estimation policy remains linear.
@@
Note that, in networked estimation, feedback can serve as a mechanism that helps the encoder know the belief of the decoder about the states of the source. Previous works on networked estimation have mostly assumed that feedback is perfect. However, when feedback is imperfect, the knowledge of the encoder about the belief of the decoder becomes distorted. This consequently can cause a degradation in the system performance. Despite its critical importance, there has been limited research on networked estimation with unreliable feedback [32, 33, 34, 35, 36, 37]. Notably, Nourian et al. [32, 33] studied the estimation of a Gauss–Markov process over a lossy channel with erroneous feedback. The goal in [32] is to minimize the trace of error variance subject to an energy harvesting constraint, and that in [33] is to minimize a convex combination of the trace of error variance and the communication energy. The authors of these articles used the notion of information state to transform the problem into a simplified one, and accordingly derived the optimal signal-independent scheduling policy and the optimal estimation policy. Khina et al. [34] studied the sequential coding of a Gauss–Markov process with multiple sensors over an i.i.d. channel with delayed feedback, and characterized the achievable causal rate-distortion region. Furthermore, Guo et al. [35] and Li et al. [36, 37] studied the estimation of a Gauss–Markov process over a lossy channel with compromised feedback, and, for some fixed scheduling policies, derived attack policies that target packet acknowledgment signals. The above results show that erroneous, delayed, or compromised feedback in general degrades the system performance. Note that, differently from the focus of our study, feedback can also be used for increasing the channel reliability, provisioning of the channel states, or signifying the receipts of control commands. For the effects of feedback in these cases, we refer the reader to [38, 39, 40, 41, 42], to [43, 44, 45, 46], and to [47, 48, 49, 50, 51, 52, 53, 54, 55], respectively.
@@ I-B Contributions and Outline
In this article, we establish a theoretical framework for networked estimation over unreliable channels with unreliable feedback. We characterize a globally optimal coding policy profile consisting of a symmetric threshold scheduling policy for the encoder and a linear estimation policy for the decoder. We show that the encoder requires to run four recursive equations, while the decoder requires to run only one recursive equation. We complement our theoretical results with a numerical analysis, and compare the performance limits of the networked system in different communication regimes. A key observation is about the variations of the system performance with respect to the feedback error rate.
@@
We emphasise that our findings present innovative contributions to the field of goal-oriented communication [56]. Our study deviates drastically from prior works. More specifically, the studies in [2, 3, 4, 5, 6, 7, 8, 9, 10] deal with networked estimation stability. In contrast, our results focus on networked estimation performance. The studies in [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] consider scalar sources and lossless feedback. On the contrary, our results hold for multi-dimensional sources and lossy feedback. Finally, the studies in [22, 23, 24, 32, 33, 35, 36, 37] derive variance-based scheduling policies that do not take advantage of realized measurements, and the studies in [25, 26, 27, 28, 29] derive compression policies that operate in a periodic manner. Differently, our results characterize a scheduling policy that depends on realized measurements and operates in a non-periodic manner.
@@
The article is organized in five sections. The current introductory section is followed by Section II, where we formulate a causal tradeoff. We present and discuss our main theoretical result in Sections III, and provide our numerical analysis in Section IV. Finally, we conclude the article in Section V.
@@ II Problem Statement
In this section, we mathematically formulate the problem of interest. In our networked system, the state of the source is observed by a sensor. The channel, connecting the sensor to the monitor, is composed of a forward and a backward channel. The encoder, collocated with the source, and the decoder, collocated with the monitor, constitute the two distributed decision makers of the networked system. We first introduce the models of the source and the channel, and then formulate a causal tradeoff between the packet rate and the mean square error. In our models, we consider a finite time horizon $N$, and assume that the time is discretized into time slots with constant durations.
@@ II-A Source Model
The source is partially observable Gauss–Markov process, and is described by the discrete-time state and output equations
$x_{k+1}$ $=A_{k}x_{k}+w_{k}$ (1)
$y_{k}$ $=C_{k}x_{k}+v_{k}$ (2)
@@
The source model given in (1) and (2) represents a wide class of physical processes. Note that such a Gauss–Markov model has been adopted extensively in the fields of communication and control, as studying this model can provide a foundation for development of more sophisticated systems. Moreover, the time-varying nature of the model allows approximation of nonlinear systems around their nominal trajectories, and the partially observable nature of it respects the fact that in reality only a noisy version of the output can be observed. Clearly, a model that is time-invariant and/or fully observable is a special case of this model.
@@ II-B Channel Model
The communication channel is a packet-erasure channel with feedback. At each time $k$, a message containing sensory information, represented by $\check{x}_{k}$, can be sent by the encoder over the forward channel, which has one-step delay and is i.i.d with forward error rate $\lambda$, to the decoder, where an estimate of the state of the process, represented by $\hat{x}_{k}$, should be computed. Let $\delta_{k}\in\{0,1\}$ be a binary variable such that $\delta_{k}=1$ if $\check{x}_{k}$ is transmitted at time $k$, and $\delta_{k}=0$ otherwise. A transmitted message in the forward channel at time $k$ is successfully received by the decoder with probability $\lambda^{c}:=1-\lambda$. Let $\gamma_{k}\in\{0,1\}$ be a binary random variable modelling the packet loss in the forward channel such that $\gamma_{k}=0$ if a packet loss occurs in the forward channel when it is used to exchange $\check{x}_{k}$, and $\gamma_{k}=1$ otherwise. Then, $\operatorname{\mathsf{p}}(\gamma_{k}=0)=\lambda$ for $k\in\mathbb{N}_{[0,N]}$. Accordingly, the output of the forward channel at time $k+1$ is specified as
$z^{f}_{k+1}=\left\{\begin{array}[]{l l}\check{x}_{k},&\ \text{if}\ \delta_{k}=1\ \wedge\ \gamma_{k}=1,\\[3.0pt] \mathfrak{F},&\ \text{if}\ \delta_{k}=1\ \wedge\ \gamma_{k}=0,\\[3.0pt] \mathfrak{E},&\ \text{otherwise}\end{array}\right.$ (6)
for $k\in\mathbb{N}_{[0,N]}$ with $z^{f}_{0}=\mathfrak{E}$ by convention, where $\mathfrak{F}$ and $\mathfrak{E}$ represent packet loss and absence of transmission in the forward channel, respectively. At each time $k$, if the forward channel is used to exchange $\check{x}_{k}$, i.e., if $\delta_{k}=1$, a packet acknowledgment is sent back by the decoder to the encoder over the backward channel, which has zero delay and is i.i.d. with backward error rate $\rho$. A transmitted message in the backward channel at time $k$ is successfully received by the encoder with probability $\rho^{c}:=1-\rho$. Let $\theta_{k}\in\{0,1\}$ be a binary random variable modelling the packet loss in the backward channel such that $\theta_{k}=0$ if a packet loss occurs in the backward channel when it is used to exchange $\gamma_{k}$, and $\theta_{k}=1$ otherwise. Then, $\operatorname{\mathsf{p}}(\theta_{k}=0)=\rho$ for $k\in\mathbb{N}_{[0,N]}$. Accordingly, the output of the backward channel at time $k+1$ is specified as
$z^{b}_{k+1}=\left\{\begin{array}[]{l l}\gamma_{k},&\ \text{if}\ \delta_{k}=1\ \wedge\ \theta_{k}=1,\\[3.0pt] \mathfrak{H},&\ \text{if}\ \delta_{k}=1\ \wedge\ \theta_{k}=0,\\[3.0pt] \mathfrak{G},&\ \text{otherwise}\end{array}\right.$ (10)
@@
    Quantization error is negligible, i.e., in a successful transmission, real-value sensory information can be conveyed from the encoder to the decoder without any error.
@@
The channel model given in (6) and (10) lends itself to the representation of the different communication regimes. In particular, we have reliable communication when $\lambda=0$ and $\rho=1$, have unreliable communication with reliable acknowledgments when $\lambda>0$ and $\rho=1$, have unreliable communication without acknowledgments when $\lambda>0$ and $\rho=0$, and have unreliable communication with unreliable acknowledgments when $\lambda>0$ and $\rho>0$. Note that the forward and backward channels operate with one-step delay and zero delay, respectively, as the main communication load is in the forward channel. Moreover, the effect of the quantization error is negligible because the number of bits in each data packet is assumed to be sufficiently large, which is a common assumption in the literature of networked estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
@@
According to (6), aside from the sensory information that can be encoded explicitly at time $k$ as $\check{x}_{k}$, implicit information can be encoded through $\mathfrak{F}$ and $\mathfrak{E}$. To illustrate this, suppose that $\delta_{k}=1$ if and only if $y_{k}\in\mathcal{A}_{k}$, where $\mathcal{A}_{k}$ can be a measurable set. Then, the decoder can infer that $y_{k}\in\mathcal{A}_{k}$ or $y_{k}\notin\mathcal{A}_{k}$ when $\mathfrak{F}$ or $\mathfrak{E}$ is detected at time $k+1$, respectively. Besides, according to (10), $\mathfrak{H}$ is detected by the encoder at time $k+1$ when a packet loss occurs in the backward channel at time $k+1$, and $\mathfrak{G}$ is detected at time $k+1$ when nothing is transmitted in the forward channel at time $k$.

### v2

# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Strategies and Limits

@@ Abstract
In this article, we establish a comprehensive theoretical framework for remote estimation in a networked system composed of a source that is observed by a sensor, a remote monitor that needs to estimate the state of the source in real time, and a communication channel that connects the source to the monitor. The source is a partially observable dynamical process, and the communication channel is a packet-erasure channel with feedback. We consider a novel communication model that captures implicit information. Our main objective is to identify the optimal strategies and the fundamental performance limits of the underlying system in the sense of a causal tradeoff between the packet rate and the mean square error when both forward and backward channels are unreliable. We characterise an optimal coding policy profile consisting of a scheduling policy for an encoder and an estimation policy for a decoder, collocated with the source and the monitor, respectively. We derive the recursive equations that must be solved online by the encoder and the decoder. In addition, we prove that the value function, originally defined over an expanding information set, admits a lower-dimensional representation depending only on two variables. We discuss the structural properties of the optimal policies, and analyse the computational complexity of an algorithm proposed for their computation. We then examine a range of special cases derived from our main theoretical results. We complement the theoretical results with a numerical analysis, and compare the performance of different remote estimation tasks in various operating regimes.
@@ I Introduction
Cyber-physical systems are intricate real-time systems that integrate digital and physical components [1]. These systems harness computation, communication, and control to optimise efficiency, adaptability, and automation in diverse domains, including smart factories, smart healthcare, and smart transportation. A significant challenge within cyber-physical systems arises from their inherent dynamism, which requires continuous status updating to capture the latest information about the physical environment [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As a step towards addressing this challenge, in this article, we develop a framework for remote state estimation in a networked system [^1] Throughout our study, a networked system is defined as a composed system operating over a communication network, where components interact via exchanged messages subject to communication constraints. composed of a source that is observed by a sensor, a remote monitor that needs to estimate the state of the source in real time, and a communication channel that connects the source to the monitor. The source is a partially observable dynamical process, and the communication channel is a packet-erasure channel with feedback. This framework is motivated by cyber-physical systems operating over packet-switching networks, where the quality of decision making tightly depends on the availability of sensory observations transmitted as data packets, while packet loss appears as an inevitable phenomenon due to the real-time nature of the operation. Our main objective is to identify the optimal strategies and the fundamental performance limits of the underlying system in the sense of a causal tradeoff between the packet rate and the mean square error when both forward and backward channels are unreliable, a setting that is crucial for designing robust decision-making architectures in modern communication-limited cyber-physical systems. To that end, we search for an optimal coding policy profile that determines the policies of an encoder [^2] In our study, the encoder is an entity that decides when and what information to transmit based on sensory information, network conditions, and system dynamics. and a decoder [^3] In our study, the decoder is an entity responsible for estimating the state of the source in real time based on received information, network conditions, and system dynamics., collocated with the source and the monitor, respectively.
@@ I-A Related Work
It is known that reliable communication close to the capacity limit can be attained with error correction subject to infinite delay or with persistent retransmission based on feedback. In remote estimation, where data is real-time, any delay more than a certain threshold is typically intolerable. Moreover, retransmission of stale information is unfavourable when fresh information can be transmitted with the same success rate instead. For these reasons, remote estimation is often conducted in an unreliable communication regime. Previous research has already recognised the severe effects of packet loss, as a consequence of communication unreliability, on stability of Kalman filtering [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In a seminal work, Sinopoli et al. [12] studied mean-square stability of Kalman filtering over an independent and identically distributed (i.i.d.) channel, and proved that there exists a critical point for the packet error rate above which the expected estimation error covariance is unbounded. To capture the temporal correlation of wireless channels, Wu et al. [16] addressed peak-covariance stability of Kalman filtering over a Gilbert–Elliott channel, and proved that there exists a critical region defined by the recovery and failure rates outside which the expected prediction error covariance is unbounded. In addition, to take into account the time variation of the strengths of wireless channels, Quevedo et al. [19] investigated mean-square stability of Kalman filtering over a fading channel with correlated gains, and established a sufficient condition that ensures the exponential boundedness of the expected estimation error covariance. Differently, Gupta et al. [21] investigated the estimation of a Gauss–Markov process over a packet-erasure channel, and showed that transmitting the minimum mean-square-error state estimate at the encoder at each time leads to the maximal information set for the decoder. The authors also obtained a necessary and sufficient condition for the packet-loss probability of an i.i.d. packet-erasure channel that guarantees the boundedness of the expected estimation error covariance. Dey et al. [22] also investigated the estimation of a scalar Gauss–Markov process under additive quantisation noise with and without channel feedback. They showed that, when feedback exists, the optimal strategy among all possible linear encoders corresponds to the transmission of the Kalman-filter innovation, and when feedback does not exist, there is a critical packet loss probability above which it is better to transmit the state rather than the innovation. Note that the above works assume that the outputs of the source are transmitted in a periodic manner.
@@
To determine the performance limits of remote estimation, however, one needs to navigate a tradeoff between communication and estimation costs. Various formulations of such a tradeoff have been explored in the literature [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 3, 34, 35, 36, 37, 38, 39, 40, 41]. In particular, the tradeoff between the packet rate and the mean square error, in which a signal-dependent scheduling policy and an estimation policy need to be found, was considered in [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 3, 42, 43]. In this formulation, which is of main interest to our study, Lipsa and Martins [24, 25] analysed the estimation of a scalar Gauss–Markov process over a lossless channel and an i.i.d. channel with reliable feedback. Using majorisation theory, they proved that the optimal scheduling policy is symmetric threshold-based, while the optimal estimation policy is linear switching. Nayyar et al.[26] extended [24] to an energy-harvesting sensor, where available energy is governed by an external stochastic process, while primarily focusing on a scalar finite-state Markov source and a lossless channel. They proved that the optimal scheduling policy is energy-dependent threshold-based and the optimal estimation policy is the one using the most recently received source state as the current estimate. They also discussed extensions to an orthogonal multi-dimensional Gauss–Markov source, in which state variables evolve independently. In this case, they showed that the optimal estimation policy becomes linear switching. Molin and Hirche [27] investigated the estimation of a scalar Markov process with symmetric noise distribution over a lossless channel, and devised an iterative algorithm that produces policies similar to those found in [24]. Chakravorty and Mahajan [28, 29, 30] studied the estimation of a scalar autoregressive Markov process with symmetric noise distribution over a lossless channel, an i.i.d. channel with reliable feedback, as well as a Gilbert–Elliott channel with reliable feedback, and proved based on renewal theory that the optimal scheduling policy is symmetric threshold-based and the optimal estimation policy is linear switching. Guo and Kostina [32] also addressed the estimation of the scalar Ornstein–Uhlenbeck process over a lossless channel, and based on majorisation theory and real induction proved that the optimal scheduling policy and the optimal estimation policy adhere to the same structures.
@@
Note that, in remote estimation, feedback serves as a mechanism that enables the encoder to gain insight into the decoder’s belief regarding the state of the source. Most existing studies on remote estimation rely on the assumption of perfect feedback. However, when feedback is imperfect, the knowledge of the encoder about the decoder’s belief becomes distorted. This consequently can cause a degradation in the system performance. Despite its crucial significance, there has been limited research on the effects of unreliable feedback on remote estimation [44, 45, 46, 47, 48, 49, 50]. Nourian et al. [44, 45] studied the estimation of a Gauss–Markov process over a lossy channel with erroneous feedback. The goal in [44] is to minimise the trace of error variance subject to an energy harvesting constraint, and that in [45] is to minimise a convex combination of the trace of error variance and the communication energy. The authors of these articles used the notion of information state to transform the problem into a simplified one, and accordingly derived the optimal signal-independent scheduling policy and the optimal estimation policy. Khina et al. [46] studied the sequential coding of a Gauss–Markov process with multiple sensors over an i.i.d. channel with delayed feedback, and characterised the achievable causal rate-distortion region. Guo et al. [47] and Li et al. [48, 49] studied the estimation of a Gauss–Markov process over a lossy channel with compromised feedback, and, for some fixed scheduling policies, derived attack policies that target packet acknowledgement signals. Vilni et al. [50] also examined status update of a finite-state Markov source when both forward and feedback channels are unreliable and the sensor is energy-harvesting, and developed policies minimising the long-term time average of a distortion function subject to an energy constraint. The above results show that erroneous, delayed, or compromised feedback in general degrades the system performance. Note that, in contrast to the focus of our study, feedback can also be used to enhance channel reliability, provide channel state information, or signal the receipt of control commands. For the effects of unreliable feedback in these cases, we refer the reader to [51, 52, 53, 54, 55], to [56, 57, 58, 59], and to [60, 61, 62, 63, 64, 65, 66, 67, 68], respectively.
Figure 1: A schematic view of remote state estimation over an unreliable channel with unreliable feedback.
@@ I-B Contributions and Outline
In this article, we establish a comprehensive theoretical framework for remote estimation over unreliable channels with unreliable feedback, by considering a novel communication model that captures implicit information. Within the established framework, we characterise a globally optimal coding policy profile consisting of a symmetric threshold-based scheduling policy for the encoder and a linear switching estimation policy for the decoder. We derive the recursive equations that must be solved online by the encoder and the decoder. In addition, we prove that the value function, originally defined over an expanding information set, admits a lower-dimensional representation depending only on two variables. We discuss the structural properties of the optimal policies, and analyse the computational complexity of an algorithm proposed for their computation. We then examine a range of special cases derived from our main theoretical results. We complement the theoretical results with a numerical analysis, and compare the performance of different remote estimation tasks in various operating regimes. A key observation is about the sensitivity of the system performance to the reliability of feedback: we illustrate that the lower the backward error rate, the better the system performance.
We emphasise that our findings contribute broadly to the emerging field of goal-oriented communication [69], which represents a paradigm shift from traditional communication designs toward designs that prioritise the transmission of information based on its relevance to specific tasks or objectives. Unlike traditional approaches that aim to transmit all data as accurately as possible, goal-oriented communication seeks to identify and convey only the most contextually significant portions of data—those that directly impact system performance. In our work, we focus on a key application of this paradigm, i.e., remote estimation in the context of cyber-physical systems. We examine how the importance of sensory data evolves over time and develop mechanisms to assess and prioritise this information sequentially. Through this study, we aim to reduce communication overhead while preserving the effectiveness of remote estimation tasks.
Note that unreliability in the forward and backward channels introduces fundamental challenges in the analysis and design of optimal strategies, which we have rigorously addressed in our paper. Technically speaking, unreliability in the forward channel leads to two forms of implicit information: events related to packet loss and absence of transmission, which provide an opportunity for the monitor to infer additional information beyond what is conveyed through the successful delivery of explicitly transmitted information. Moreover, unreliability in the backward channel compels the encoder to estimate the belief of the decoder, whose task is to estimate the state of the source, at each time. Together, these challenges underscore why designing optimal strategies requires careful structural characterisation.
@@
Our study departs significantly from prior work in several aspects. More specifically, the studies in [12, 13, 14, 15, 16, 17, 18, 19, 20] primarily address stability of remote estimation. In contrast, our focus is on performance optimisation. The studies in [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 3] mainly focus on scalar sources, lossless channels, or channels with lossless feedback. On the contrary, our results hold for multi-dimensional sources and unreliable channels with unreliable feedback. The studies in [50, 4, 5] focus on finite-state sources, while our work deals with continuous-state sources. Finally, the studies in [34, 35, 36, 44, 45, 47, 48, 49] derive variance-based scheduling policies that do not take advantage of realised measurements, and the studies in [37, 38, 39, 40, 41, 46, 21, 22] derive coding policies that operate in a periodic manner. Differently, our results characterise a scheduling policy that depends on realised measurements and operates in a non-periodic manner based on the value of information.
@@
The article is organised in eight sections. This introductory section is followed by Section II, where we formulate the tradeoff problem. We present and discuss our main theoretical results in Sections III, IV, and V. We explore several special cases that emerge from our main theoretical results in Section VI. We provide our numerical analysis in Section VII. Finally, we conclude the article in Section VIII.
I-C Preliminaries
In the sequel, the sets of real numbers and non-negative integers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. For $x,y\in\mathbb{N}$ and $x\leq y$, the set $\mathbb{N}_{[x,y]}$ denotes $\{z\in\mathbb{N}|x\leq z\leq y\}$. The sequence of all vectors $x_{t}$, $t=p,\dots,q$, is represented by $\boldsymbol{x}_{p:q}$. For matrices $X$ and $Y$, the relations $X\succ 0$ and $Y\succeq 0$ denote that $X$ and $Y$ are positive definite and positive semi-definite, respectively. The indicator function of a subset $\mathcal{A}$ of a set $\mathcal{X}$ is denoted by $\mathds{1}_{\mathcal{A}}:\mathcal{X}\to\{0,1\}$. The symbol $\mathcal{O}$ describes the upper bound on the growth rate of a function. The symmetric decreasing rearrangement of a measurable function $f(x)$ vanishing at infinity is represented by $f^{*}(x)$. The probability measure of a random variable $x$ is represented by $\mathsf{P}(x)$, its probability density or probability mass function by $\operatorname{\mathsf{p}}(x)$, and its expected value and covariance by $\operatorname{\mathsf{E}}[x]$ and $\operatorname{\mathsf{cov}}[x]$, respectively. Stochastic kernels will be adopted to represent decision policies. Let $(\mathcal{X},\mathcal{B}_{\mathcal{X}})$ and $(\mathcal{Y},\mathcal{B}_{\mathcal{Y}})$ be two measurable spaces. A measurable stochastic kernel $\operatorname{\mathsf{P}}:\mathcal{B}_{\mathcal{Y}}\times\mathcal{X}\to[0,1]$ is a mapping such that $\mathcal{A}\mapsto\operatorname{\mathsf{P}}(\mathcal{A}|x)$ is a probability measure on $(\mathcal{Y},\mathcal{B}_{\mathcal{Y}})$ for any $x\in\mathcal{X}$, and $x\mapsto\operatorname{\mathsf{P}}(\mathcal{A}|x)$ is a measurable function for any $\mathcal{A}\in\mathcal{B}_{\mathcal{Y}}$.
@@ II Problem Statement
In this section, we mathematically formulate the problem of interest. Our networked system is illustrated in Fig. 1. The state of the source is observed by the sensor. The channel, connecting the sensor to the monitor, is composed of a forward and a backward channel. The encoder, collocated with the source, and the decoder, collocated with the monitor, constitute the two distributed decision makers of the networked system. We first introduce the models of the source and the channel, and then formulate a causal tradeoff between the packet rate and the mean square error. Throughout this study, we consider a finite time horizon, denoted by $N$, and assume that the time is discretised into time slots with constant durations.
@@ II-A Source Model
The source is a partially observable Gauss–Markov process, and is described by the discrete-time state and output equations
$x_{k+1}$ $=A_{k}x_{k}+w_{k},$ (1)
$y_{k}$ $=C_{k}x_{k}+v_{k},$ (2)
@@
The source model given in (1) and (2) represents a wide class of dynamical processes. Note that such a Gauss–Markov model has been adopted extensively in the fields of communication and control, since this model provides a foundation for development of more sophisticated systems. Moreover, the time-varying nature of the model allows approximation of nonlinear systems around their nominal trajectories, and its partially observable nature respects the fact that in reality only a noisy version of the output can be observed. Clearly, a model that is time-invariant and/or fully observable is a special case of this model.
@@ II-B Channel Model
Both forward and backward channels are packet-erasure channels. At each time $k$, a message containing sensory information, represented by $\check{x}_{k}$, can be sent by the encoder over the forward channel, which has one-step delay and is i.i.d. with forward error rate $\lambda$, to the decoder, where an estimate of the state of the process, represented by $\hat{x}_{k}$, should be computed. Let $u_{k}\in\{0,1\}$ be a binary variable such that $u_{k}=1$ if $\check{x}_{k}$ is transmitted at time $k$, and $u_{k}=0$ otherwise. A transmitted message in the forward channel at each time is successfully received by the decoder with probability $\lambda^{c}:=1-\lambda$. Let $\gamma_{k}\in\{0,1\}$ be a binary random variable modelling the packet loss in the forward channel such that $\gamma_{k}=0$ if a packet loss occurs in the forward channel when it is used to exchange $\check{x}_{k}$, and $\gamma_{k}=1$ otherwise. Then, $\operatorname{\mathsf{p}}(\gamma_{k}=0)=\lambda$ for $k\in\mathbb{N}_{[0,N]}$. Accordingly, the output of the forward channel at time $k+1$ is specified as
$z^{f}_{k+1}=\left\{\begin{array}[]{l l}\check{x}_{k},&\ \text{if}\ u_{k}=1\ \wedge\ \gamma_{k}=1,\\[3.0pt] \mathfrak{F},&\ \text{if}\ u_{k}=1\ \wedge\ \gamma_{k}=0,\\[3.0pt] \mathfrak{E},&\ \text{otherwise},\end{array}\right$. (6)
for $k\in\mathbb{N}_{[0,N]}$ with $z^{f}_{0}=\mathfrak{E}$ by convention, where $\mathfrak{F}$ and $\mathfrak{E}$ represent packet loss and absence of transmission in the forward channel, respectively. At each time $k$, if the forward channel is used to exchange $\check{x}_{k}$, i.e., if $u_{k}=1$, a packet acknowledgement is sent back by the decoder to the encoder over the backward channel, which has zero delay and is i.i.d. with backward error rate $\rho$. A transmitted message in the backward channel at each time is successfully received by the encoder with probability $\rho^{c}:=1-\rho$. Let $\theta_{k}\in\{0,1\}$ be a binary random variable modelling the packet loss in the backward channel such that $\theta_{k}=0$ if a packet loss occurs in the backward channel when it is used to exchange $\gamma_{k}$, and $\theta_{k}=1$ otherwise. Then, $\operatorname{\mathsf{p}}(\theta_{k}=0)=\rho$ for $k\in\mathbb{N}_{[0,N]}$. Accordingly, the output of the backward channel at time $k+1$ is specified as
$z^{b}_{k+1}=\left\{\begin{array}[]{l l}\gamma_{k},&\ \text{if}\ u_{k}=1\ \wedge\ \theta_{k}=1,\\[3.0pt] \mathfrak{H},&\ \text{if}\ u_{k}=1\ \wedge\ \theta_{k}=0,\\[3.0pt] \mathfrak{G},&\ \text{otherwise},\end{array}\right$. (10)
@@
    By convention, set $\gamma_{k}=\theta_{k}=0$ when nothing is transmitted in the forward channel at time $k$, i.e., $u_{k}=0$.
  3. (iii)
    Quantisation error is negligible, i.e., in a successful transmission, real-value sensory information can be conveyed from the encoder to the decoder without any error.
@@
The channel model given in (6) and (10) lends itself to the representation of different communication regimes. In particular, we have reliable communication [^4] Throughout this study, reliable communication and reliable feedback refer to the idealised cases $\lambda=0$ and $\rho=0$, respectively. We use these regimes as theoretical benchmarks. In practice, reliability typically implies very low but non-perfect error rates (e.g., <0.0001<0.0001). without acknowledgements when $\lambda=0$ and $\rho=1$, unreliable communication with reliable acknowledgements when $\lambda>0$ and $\rho=0$, unreliable communication without acknowledgements when $\lambda>0$ and $\rho=1$, and unreliable communication with unreliable acknowledgements when $\lambda>0$ and $\rho>0$ (i.e., the general regime). Note that the forward and backward channels are assumed to operate with one-step delay and zero delay, respectively, as the main communication load is in the forward channel. Moreover, the effect of the quantisation error is assumed to be negligible, as the number of bits in each data packet is often sufficiently large in packet-switching networks, which is a common assumption in the literature of remote estimation [23, 24, 25, 27, 28, 29, 30, 31, 32, 33, 3, 34, 35, 36].
@@
Note that according to (6), aside from the sensory information that can be encoded explicitly at time $k$ as $\check{x}_{k}$, implicit information can also be encoded through $\mathfrak{F}$ and $\mathfrak{E}$. To illustrate this, suppose that $u_{k}=1$ if and only if $y_{k}\in\mathcal{A}_{k}$, where $\mathcal{A}_{k}$ can be any measurable set. Then, the decoder can infer that $y_{k}\in\mathcal{A}_{k}$ or $y_{k}\notin\mathcal{A}_{k}$ when $\mathfrak{F}$ or $\mathfrak{E}$ is detected at time $k+1$, respectively. Moreover, note that according to (10), $\mathfrak{H}$ is detected by the encoder at time $k+1$ when sensory information is transmitted in the forward channel at time $k$ and a packet loss occurs in the backward channel at time $k+1$, and $\mathfrak{G}$ is detected at time $k+1$ when nothing is transmitted in the backward channel at time $k$.